# Modeling the TMD shape function in $J/\psi$ electroproduction

### v2

@@ 3.2 Effect on the cross-section
where we use the $C_{g/a}$ at NLO, see e.g., the Ref. Echevarria:2015uaa . Moreover, the evolved TMDPDF at an arbitrary scales $(\mu_{f},\zeta_{f})$ is obtained from its value at the initial scales $(\mu_{i},\zeta_{i})$, using the evolution kernel:
@@
The function $\gamma_{V}$ refers to the finite part of the renormalization of the vector form factor, which is known at N3LO Lee:2022nhh .
@@
The specific values of $B_{f}(x)$ depending on $x$ are shown in Ref. Bor:2022fga . Also, $\mu_{\text{NP}}$ sets the separation between the perturbative and the NP behavior of the TMDPDF.
@@
In Fig. 4, we illustrate the impact of the TMDShF on the TMD evolution of the convolution. [^4] It should be noted that a superscript $g$ has been introduced to indicate the evolution of the TMD gluon distribution, while a superscript $S$ denotes the evolution of the TMDShF. Specifically, we show $e^{-S_{A}^{g}}$ at NNLL accuracy together with the third lane of Eq. (25), evaluated with $\gamma_{s}$ at NLO. The uncertainty bands reflect the variation of the parameter $\zeta_{B,b}$ by a factor of two around its default value, $\zeta_{B,b}=1$. Since the evolution depends on the $b_{T}^{*}$-prescription, and there is no unique choice for $b_{T\text{max}}$, we present results for three representative values. The choice $b_{T\text{max}}=0.5$ GeV-1 is clearly too small for $\mu_{H}=3,6$ GeV, as the distribution falls below unity already for $b_{T}\leq b_{T,\text{min}}$. Conversely, $b_{T\text{max}}=2$ GeV-1 lies beyond the region where the strong coupling remains perturbative; nevertheless, it is instructive to explore at which point the exponential factor for $\mu_{H}=3$ GeV starts decreasing within the perturbative domain. Overall, we find that the effect of the TMDShF on the Sudakov factor of the convolution is to shift the distribution upward and toward larger $b_{T}$ values. This occurs because $\gamma_{s}<0$, thereby flipping the sign of the exponential. Furthermore, the relative deviation associated with the theoretical uncertainty from $\zeta_{B,b}$ exhibits a universal behavior across energies and $b_{T\text{max}}$: it grows rapidly at small $b_{T}$ and saturates at maximum values of 2.8, 6.8 and 24.3% for $b_{T\text{max}}=$ 0.5, 1 and 2 GeV-1, respectively.
@@
To explore the impact of the TMDShF in the NP region, we show in Fig. 5 the quantity $q_{T}\cdot\mathcal{C}[f_{1}^{g}S_{,^{1}S_{0}^{[8]}}]$, interpreted as a probability distribution, at the highest level of accuracy considered. Results are presented for $A_{S}=B_{S}=0$ (dashed curves) and for $A_{S}=0.16$ GeV2, $B_{S}=0.5$ GeV2 (solid curves). The remaining parameters are fixed to $A_{f}=0.16$ GeV2, $\mu_{\text{NP}}=1$ GeV, $\zeta_{B,0}=0.1$, and $\zeta_{B,b}=0.5$. For the parton distribution functions and LDMEs we adopt the MSTW2008lo and SV sets, respectively. We restrict the plots to $q_{T}\leq\mu_{H}/2$, which we identify as the region of validity of TMD factorization. The convolution exhibits a strong dependence on the gluon PDF, and we therefore display results for several values of $x_{A}$. The characteristic behavior of the gluon PDF is observed: for $x_{A}=0.1$, the peak of the distribution grows as the energy decreases. At lower values of $x_{A}$, however, the peaks for $\mu_{H}=3$ and $6$ GeV approach those of higher energies, reflecting the faster fall-off of the gluon PDF at small $x_{A}$ and low scales—most prominently for $x_{A}=10^{-3}$. Another general consequence of gluon PDF behavior is that, for any fixed $\mu_{H}$, the peak height decreases with increasing $x_{A}$. Finally, the uncertainty bands represent the effect of the TMDShF in the non-perturbative region, evaluated for $(A_{S},B_{S})=(0.16,0.1)$ GeV2, corresponding to values discussed in the previous section. As expected, the width of the bands decreases with increasing $\mu_{H}$, since the exponential suppression of the NP model becomes stronger at larger hard scales.
@@
where the superscript $\ell p$ highlight that this is the NP model for the leptoproduction cross-section. While the above parametrization offers a simple and practical framework for exploring the low–transverse-momentum region, where the TMDShF plays a significant role, a more refined modeling of $S_{\text{NP}}$—or of the $b_{T}$-prescription employed to separate the perturbative from the non-perturbative domain—will ultimately be required for precision phenomenology. For the present study, and in line with global fits to low-energy SIDIS data as well as high-energy Drell–Yan and $Z$-boson production data, we adopt $A=0.414$ GeV2 and $\tilde{\mu}_{\text{NP}}=1.6$ GeV as central values, and vary $A$ within the range discussed above when presenting the predictions of the next section.
@@ 4 Theoretical predictions
QED coupling const. $\alpha_{em}$ $\alpha_{em}=137$
QCD coupling const. $\alpha_{s}\equiv 4\pi\,a_{s}$ $\alpha_{s}(M)\simeq 0.261$ taken at three-loop ParticleDataGroup:2024cfk
WW function $f_{\gamma/e}$ $Q_{min}^{2}=m_{e}^{2}y^{2}/(1-y),Q_{max}^{2}=2.5GeV^{2}$ Frixione:1993yw
@@
Cusp AD $\Gamma_{cusp}$ $\mathcal{O}(a_{s}^{3})$ Vladimirov:2017ksc
TMDPDF AD $\gamma_{V}$ $\mathcal{O}(a_{s}^{2})$ Echevarria:2015uaa
TMDShF AD $\gamma_{s}$ $\mathcal{O}(a_{s}^{1})$ Echevarria:2024idp
@@
Collins-Soper kernel $\mathcal{D}$
Pert. part $\mathcal{D}_{pert}$ $\mathcal{O}(a_{s}^{2})$ Echevarria:2012pw
@@
Unpolarized gluon TMDPDF $f_{1}^{g}$
Matching coef. $C_{f\leftarrow f^{\prime}}$ $\mathcal{O}(\alpha_{s}^{1}$) Echevarria:2015uaa
Unpol. PDF $f_{1}$ NNLO from MSTW2008 PDF set Martin_2009
@@
Unpolarized $[n]$-TMDShF $S_{[n]}$
Matching coef. $C_{[n]}^{[m]}$ $\mathcal{O}(a_{s}^{1})$ Echevarria:2024idp
[n]LDME $\left<\mathcal{O}_{[n]}\right$> We use several sets (see in the text) brambilla2024 ; Chao_2012 ; Sharma_2013 ; Bodwin_2014
@@
Our predictions are obtained within the TMD factorization framework, working at leading power in the $P_{\psi\perp}/Q$ expansion. Large logarithms are resummed to NNLL accuracy, and the perturbative hard function is included at $\mathcal{O}(\alpha_{s})$ with full one-loop virtual corrections. The analysis is restricted to the color-octet channel, which is expected to dominate in the inelastic regime at large photon virtuality, $Q^{2}\gg M^{2}$. By contrast, the color-singlet contribution appears only at $\mathcal{O}(\alpha_{s}^{2})$ and is suppressed in the strong coupling, while diffractive contributions at $z=1$ are further power suppressed by $1/Q^{2}$ relative to the CO channel Fleming:1997fq .
@@
For the theoretical estimates, we choose two extreme center-of-mass energy, $\sqrt{S}=140$ and $45$ GeV. We show the $P_{\psi\perp}$ spectrum for $P_{\psi\perp}<\mu_{H}/2$, a reliable TMD factorization domain. We used MSTW2008 at NNLO PDF set Martin:2009iq .
In Fig. 6, we present the $P_{\psi\perp}$-spectrum of the differential cross-section for $\sqrt{s}=140\,\text{GeV}$ (a,b) and $\sqrt{s}=45\,\text{GeV}$ (c,d) for different values of $x_{B}$ and $Q=4,9$ GeV. We observe a peaked structure in the spectrum, with the position of the peak shifting toward larger $P_{\psi\perp}$ and becoming narrower as $Q$ increases which is consistent with expectations from resummed TMD calculations as the behavior arises from stronger Sudakov suppression at higher scales, and from the width of the NP effect. Moreover, we observe a significant change in the profile of the distribution in the TMD region when comparing the plots for $Q=4$ GeV and $Q=9$ GeV. This behavior is attributed to the fact that the hard scale $\mu_{H}$ increases quadratically with $Q$, and consequently, the condition $P_{\psi\perp}<\mu_{H}/2$, under which our prediction is reliable, covers almost the entire $P_{\psi\perp}$-spectrum where the cross-section is not zero. This result may indicate that the energy $Q=9$ GeV is too large for our factorization framework, potentially requiring (at this energy and above), a proper scale separation between $Q$ and $M$. In other words, a specific TMD factorization approach valid for $Q\gg M$ may be necessary. We leave this analysis to future work. To conclude with this figure, we include uncertainty bands arising from variations of the rapidity and renormalization scales by a factor of two. Although not explicitly shown here, we observe narrower uncertainty bands at higher resummation accuracy, as expected.
@@
Figure 8: Differential cross-section as functions of $J/\psi$ transverse momentum $d\sigma_{ep}/dP_{\psi\perp}$ for different sets of LDMEs: BBXWbrambilla2024 , CMSWZChao_2012 , SVSharma_2013 , and BCKLBodwin_2014 .
@@
The final analysis in our predictions comes due to the different LDME sets. In Fig. 6 and 7, we show the prediction using SV LDME set Sharma_2013 . However, we have different sets of LDMEs in the literature, where some of the sets are extracted from global fit, but with high $P_{T}$ data. In particular, each line in Fig. 8 shows the NNLL prediction obtained using the following LDME sets: BBXW brambilla2024 , CMSWZ Chao_2012 , SV and BCKL Bodwin_2014 We see a wider spectrum using different LDMEs than in the theoretical uncertainty band, which tells a better fit is needed to extract them, especially in low-$P_{T}$.
@@ 4.1 On the TMDShF influence
which directly quantifies the suppression induced by the TMD shape function in the low-$P_{\psi\perp}$ region. Here $d{\sigma}_{f_{1}^{g}\otimes S_{\psi}}$ and $d\sigma_{f_{1}^{g}\otimes<O_{J/\psi}^{[m]}>}$ denote the differential cross-section with and without TMDShF, respectively. The results are displayed in Fig. 9 for EIC kinematics, considering $B_{S}=0,$ 1.5 and 3 GeV2 and varying $\mu_{H}$, with the spectrum plotted up to $P_{\psi\perp}=\mu_{H}/2$. We observe that, unlike the case $B_{S}\neq 0$, the prediction including the TMDShF overshoots the standard NRQCD result. For the left panel, where the NP contribution of the TMDShF is effectively switched off, the ratio remains close to unity, and the residual variation (within about 20%) originates mainly from scale dependence. As $B_{S}$ increases, corresponding to an enhanced strength of the TMDShF, a clear distortion of the ratio develops. For $B_{S}=1.5$ GeV2 exhibits a pronounced peak around $P_{\psi\perp}\sim 5-6$ GeV. This indicates a moderate enhancement of the cross-section in this intermediate momentum region, likely arising from constructive interference or modified phase-space weighting due to the inclusion of the TMDShF. The dependence on the scale becomes stronger in this regime. For $B_{S}=3$ GeV2, the effect becomes even more pronounced. The peak in the intermediate region increases, while at larger momenta the ratio falls below unity. The transition point between enhancement and suppression is roughly stable across $\mu_{H}$, indicating a dynamical origin tied to the intrinsic shape of the additional function rather than to the perturbative scale choice. Overall, the results demonstrate that the inclusion of the TMDShF governed by $B_{S}$ introduces nontrivial shape modifications to the transverse-momentum distribution. Increasing $B_{S}$ amplifies these distortions, confirming its role as a strength or broadness parameter of the TMDShF. We emphasize, however, that this behavior pertains to the specific implementation described in Eq. (28), and could change under alternative NP parametrization of the CS kernel and the gluon TMDPDF.
@@
It is worth to emphasize that the present analysis has been confined to the case of electroproduction. Nonetheless, the theoretical framework developed in this work can be straightforwardly extended to provide predictions for photoproduction processes. It is, however, well established that the H1 measurements of large-$p_{T}$ $J/\psi$ photoproduction H1:2010udv are accurately reproduced by the color-singlet contribution alone, computed at $\mathcal{O}(\alpha_{s}^{2})$ Flore:2020jau , where the CO contribution at $\mathcal{O}(\alpha_{s})$ is known to significantly overshoot the experimental data. Moreover, theoretical predictions based on the conventional NRQCD factorization framework are known to fail in describing the data, particularly in the low-$p_{T}$ and high-$z$ region, as discussed in Ref. Brambilla:2024iqg . In a preliminary analysis, as a qualitative comparison, we confront our estimate for photoproduction at $z\sim 1$ with the H1 measurement in the $0.75<z<0.9$ bin H1:2010udv . We find that the incorporation of the TMDShF within the TMD framework leads to theoretical predictions of the CO channel that exhibit markedly better agreement with the experimental data, yielding lower cross sections at small $p_{T}$ relative to the current collinear NRQCD results. However, we leave a detailed analysis of photoproduction in the TMD framework for future work.

### v3

@@ 3.2 Effect on the cross-section
where we use the $C_{g/a}$ at NLO, see e.g., the Ref. Echevarria et al. (2015). Moreover, the evolved TMDPDF at an arbitrary scales $(\mu_{f},\zeta_{f})$ is obtained from its value at the initial scales $(\mu_{i},\zeta_{i})$, using the evolution kernel:
@@
The function $\gamma_{V}$ refers to the finite part of the renormalization of the vector form factor, which is known at N3LO Lee et al. (2022).
@@
The specific values of $B_{f}(x)$ depending on $x$ are shown in Ref. Bor and Boer (2022). Also, $\mu_{\text{NP}}$ sets the separation between the perturbative and the NP behavior of the TMDPDF.
@@
In Fig. 4, we illustrate the impact of the TMDShF on the TMD evolution of the convolution. [^4] It should be noted that a superscript $g$ has been introduced to indicate the evolution of the TMD gluon distribution, while a superscript $S$ denotes the evolution of the TMDShF. Specifically, we show $e^{-S_{A}^{g}}$ at NNLL accuracy together with the third lane of Eq. (25), evaluated with $\gamma_{s}$ at NLO. The uncertainty bands reflect the variation of the parameter $\zeta_{B,b}$ by a factor of two around its default value, $\zeta_{B,b}=1$. Since the evolution depends on the $b_{T}^{*}$-prescription, and there is no unique choice for $b_{T\text{max}}$, we present results for three representative values. The choice $b_{T\text{max}}=0.5$ GeV-1 is clearly too small for $\mu_{H}=3,6$ GeV, as the distribution falls below unity already for $b_{T}\leq b_{T,\text{min}}$. Conversely, $b_{T\text{max}}=2$ GeV-1 lies beyond the region where the strong coupling remains perturbative; nevertheless, it is instructive to explore at which point the exponential factor for $\mu_{H}=3$ GeV starts decreasing within the perturbative domain. Overall, we find that the effect of the TMDShF on the Sudakov factor of the convolution is to shift the distribution upward and toward larger $b_{T}$ values. This occurs because $\gamma_{s}<0$, thereby flipping the sign of the exponential. Furthermore, the relative scale uncertainty, generally defined as $\delta=(f_{max}-f_{min})/2f_{c}$ with $f_{max/min}$ the maximum and minimum values of the envelope, and $f_{c}$ the central value, associated to a 3-point variation of $\zeta_{B,b}$ exhibits a universal behavior across energies and $b_{T\text{max}}$. It grows rapidly at small $b_{T}$ and saturates at maximum values of 2.8, 6.8 and 24.3% for $b_{T\text{max}}=$ 0.5, 1 and 2 GeV-1, respectively.
@@
To explore the impact of the TMDShF in the NP region, we show in Fig. 5 the quantity $q_{T}\cdot\mathcal{C}[f_{1}^{g}S_{{}^{1}S_{0}^{[8]}}$], interpreted as a probability distribution, at the highest level of accuracy considered. Results are presented for $A_{S}=B_{S}=0$ (dashed curves) and for $A_{S}=0.16$ GeV2, $B_{S}=0.5$ GeV2 (solid curves). The remaining parameters are fixed to $A_{f}=0.16$ GeV2, $\mu_{\text{NP}}=1$ GeV, $\zeta_{B,0}=0.1$, and $\zeta_{B,b}=0.5$. For the parton distribution functions and LDMEs we adopt the MSTW2008lo and SV sets, respectively. We restrict the plots to $q_{T}\leq\mu_{H}/2$, which we identify as the region of validity of TMD factorization. The convolution exhibits a strong dependence on the gluon PDF, and we therefore display results for several values of $x_{A}$. The characteristic behavior of the gluon PDF is observed: for $x_{A}=0.1$, the peak of the distribution grows as the energy decreases. At lower values of $x_{A}$, however, the peaks for $\mu_{H}=3$ and $6$ GeV approach those of higher energies, reflecting the faster fall-off of the gluon PDF at small $x_{A}$ and low scales—most prominently for $x_{A}=10^{-3}$. Another general consequence of gluon PDF behavior is that, for any fixed $\mu_{H}$, the peak height decreases with increasing $x_{A}$. Finally, the uncertainty bands represent the effect of the TMDShF in the non-perturbative region, evaluated for $(A_{S},B_{S})=(0.16,0.1)$ GeV2, corresponding to values discussed in the previous section. As expected, the width of the bands decreases with increasing $\mu_{H}$, since the exponential suppression of the NP model becomes stronger at larger hard scales.
@@
where the superscript $\ell p$ highlight that this is the NP model for the leptoproduction cross-section. While the above parametrization offers a simple and practical framework for exploring the low–transverse-momentum region, where the TMDShF plays a significant role, a more refined modeling of $S_{\text{NP}}$—or of the $b_{T}$-prescription employed to separate the perturbative from the non-perturbative domain—will ultimately be required for precision phenomenology. For the present study, and in line with global fits to low-energy SIDIS data as well as high-energy Drell–Yan and $Z$-boson production data, we adopt $A=0.414$ GeV2 and $\tilde{\mu}_{\text{NP}}=1.6$ GeV as central values, and vary $A$ within the range discussed above when presenting the predictions of the next section. Whereas, $B_{f}(x)$ is parameterized as
$B_{f}(x)=\frac{9}{8}(-g_{1}\log(2)+g_{2}(1+2g_{3}\log(\frac{10xx_{0}}{x_{0}-x})))$ (29)
with $g_{1}=0.184~\text{GeV}^{2}$, $g_{2}=0.201~\text{GeV}^{2}$, $g_{3}=-129~\text{GeV}^{2}$ and $x_{0}=0.009$ Aybat and Rogers (2011). A point to note is that the parameters were obtained for the case of quark TMDs (two); hence, to use them for the gluon TMD, an additional factor of $9/8(\equiv C_{A}/(2C_{F}))$ is multiplied.
@@ 4 Theoretical predictions
QED coupling const. $\alpha_{em}$ $\alpha_{em}=1/137$
QCD coupling const. $\alpha_{s}\equiv 4\pi\,a_{s}$ $\alpha_{s}(M)\simeq 0.261$ taken at three-loop Navas and others (2024)
WW function $f_{\gamma/e}$ $Q_{min}^{2}=m_{e}^{2}y^{2}/(1-y),Q_{max}^{2}=2.5GeV^{2}$ Frixione et al. (1993)
@@
Cusp AD $\Gamma_{cusp}$ $\mathcal{O}(a_{s}^{3})$ Vladimirov (2018)
TMDPDF AD $\gamma_{V}$ $\mathcal{O}(a_{s}^{2})$ Echevarria et al. (2015)
TMDShF AD $\gamma_{s}$ $\mathcal{O}(a_{s}^{1})$ Echevarria et al. (2024)
@@
[…]
Pert. part $\mathcal{D}_{pert}$ $\mathcal{O}(a_{s}^{2})$ Echevarria et al. (2013)
@@
[…]
Matching coef. $C_{f\leftarrow f^{\prime}}$ $\mathcal{O}(a_{s}^{1}$) Echevarria et al. (2015)
Unpol. PDF $f_{1}$ NNLO from MSTW2008 PDF set Martin et al. (2009a)
@@
[…]
Matching coef. $C_{[n]}^{[m]}$ $\mathcal{O}(a_{s}^{1})$ Echevarria et al. (2024)
[n]LDME $\left<\mathcal{O}_{[n]}\right$> We use several sets (see in the text) Brambilla et al. (2024b); Chao et al. (2012); Sharma and Vitev (2013); Bodwin et al. (2014)
@@
Our predictions are obtained within the TMD factorization framework, working at leading power in the $P_{\psi\perp}/Q$ expansion. Large logarithms are resummed to NNLL accuracy, and the perturbative hard function is included at $\mathcal{O}(\alpha_{s})$ with full one-loop virtual corrections. The analysis is restricted to the color-octet channel, which is expected to dominate in the inelastic regime at large photon virtuality, $Q^{2}\gg M^{2}$. By contrast, the color-singlet contribution appears only at $\mathcal{O}(\alpha_{s}^{2})$ and is suppressed in the strong coupling, while diffractive contributions at $z=1$ are further power suppressed by $1/Q^{2}$ relative to the CO channel Fleming and Mehen (1998).
@@
For the theoretical estimates, we choose two extreme center-of-mass energy, $\sqrt{S}=140$ and $45$ GeV. We show the $P_{\psi\perp}$ spectrum for $P_{\psi\perp}<\mu_{H}/2$, a reliable TMD factorization domain. We used MSTW2008 at NNLO PDF set Martin et al. (2009b). The minimum scale of the PDF is 1 GeV.
In Fig. 6, we present the $P_{\psi\perp}$-spectrum of the differential cross-section for $\sqrt{s}=140\,\text{GeV}$ (a,b) and $\sqrt{s}=45\,\text{GeV}$ (c,d) for different values of $x_{B}$ and $Q=4,9$ GeV. We observe a peaked structure in the spectrum, with the position of the peak shifting toward larger $P_{\psi\perp}$ and becoming narrower as $Q$ increases which is consistent with expectations from resummed TMD calculations as the behavior arises from stronger Sudakov suppression at higher scales, and from the width of the NP effect. Moreover, we observe a significant change in the profile of the distribution in the TMD region when comparing the plots for $Q=4$ GeV and $Q=9$ GeV. This behavior is attributed to the fact that the hard scale $\mu_{H}$ increases quadratically with $Q$, and consequently, the condition $P_{\psi\perp}<\mu_{H}/2$, under which our prediction is reliable, covers almost the entire $P_{\psi\perp}$-spectrum where the cross-section is not zero. This result may indicate that the energy $Q=9$ GeV is too large for our factorization framework, potentially requiring (at this energy and above), a proper scale separation between $Q$ and $M$. In other words, a specific TMD factorization approach valid for $Q\gg M$ may be necessary. We leave this analysis to future work. To conclude with this figure, we include uncertainty bands arising from variations of the rapidity and renormalization scales by a factor of two. Although not explicitly shown here, we observe narrower uncertainty bands at higher resummation accuracy, as expected. Moreover, one obtain the oscillatory scale uncertainty band at large $P_{\psi\perp}$ is the consequence of the minimum scale of the PDF being 1 GeV and higher.
@@
Figure 8: Differential cross-section as functions of $J/\psi$ transverse momentum $d\sigma_{ep}/dP_{\psi\perp}$ for different sets of LDMEs: BBXWBrambilla et al. (2024b), CMSWZChao et al. (2012), SVSharma and Vitev (2013), and BCKLBodwin et al. (2014).
@@
Moreover, we obtain a negative (unphysical) cross-section as $P_{\psi\perp}\to\mu_{H}/2$, particularly for large $Q(9~\text{GeV})$ at large $x$. This is due to the behavior of TMD convolution in that domain; see Fig. 5. Such behavior of a negative cross-section has also been reported in Ref Boer et al. (2023). However, with a change (increase) in the value of NP parameters, which means suppressing the non-perturbative contributions further, we can obtain the positive cross-section even at the large $P_{\psi\perp}\to\mu_{H}/2$. However, strictly speaking, the TMD factorization formalism is applicable for $P_{\psi\perp}<<\mu_{H}$. The appearance of negative values near $P_{\psi\perp}\sim\mu_{H}/2$ signals the breakdown of the TMD expansion near the (assumed) boundary of its applicability, rather than an inconsistency of the formalism within its domain of validity. Additionally, we do not have constraints on the non-perturbative parameters yet, particularly for the gluon TMD, from the experiments. This leads to the conclusion that a better parameterization is required, particularly for the parameter that depends on $x$; $B_{f}(x)$. Moreover, particularly at large $P_{\psi\perp}$, power correction and higher twist effect might play a sizable role. We left these analysis for future work.
The final analysis in our predictions comes due to the different LDME sets. In Fig. 6 and 7, we show the prediction using SV LDME set Sharma and Vitev (2013). However, we have different sets of LDMEs in the literature, where some of the sets are extracted from global fit, but with high $P_{T}$ data. In particular, each line in Fig. 8 shows the NNLL prediction obtained using the following LDME sets: BBXW Brambilla et al. (2024b), CMSWZ Chao et al. (2012), SV and BCKL Bodwin et al. (2014) We see a wider spectrum using different LDMEs than in the theoretical uncertainty band, which tells a better fit is needed to extract them, especially in low-$P_{T}$.
@@ 4.1 On the TMDShF influence
which directly quantifies the suppression induced by the TMD shape function in the low-$P_{\psi\perp}$ region. Here $d{\sigma}_{f_{1}^{g}\otimes S_{\psi}}$ and $d\sigma_{f_{1}^{g}\otimes<O_{J/\psi}^{[m]}>}$ denote the differential cross-section with and without TMDShF, respectively. The results are displayed in Fig. 9 for EIC kinematics, considering $B_{S}=0,$ 1.5 and 3 GeV2 and varying $\mu_{H}$, with the spectrum plotted up to $P_{\psi\perp}=\mu_{H}/2$. We used SV LDMEs set for these plots. Moreover, the LDME sets have a negligible or no effect on its behavior. The reason lies in the perturbative tails, as they are the same for $S$- and $P$- states except for the cross terms, which we have considered zero in our numerical estimates as discussed in section 3. Hence, the ratio mostly represents the effect of the shape function.
@@
We observe that, unlike the case $B_{S}\neq 0$, the prediction including the TMDShF overshoots the standard NRQCD result. For the left panel, where the NP contribution of the TMDShF is effectively switched off, the ratio remains close to unity, and the residual variation (within about 20%) originates mainly from scale dependence. As $B_{S}$ increases, corresponding to an enhanced strength of the TMDShF, a clear distortion of the ratio develops. For $B_{S}=1.5$ GeV2 exhibits a pronounced peak around $P_{\psi\perp}\sim 5-6$ GeV. This indicates a moderate enhancement of the cross-section in this intermediate momentum region, likely arising from constructive interference or modified phase-space weighting due to the inclusion of the TMDShF. The dependence on the scale becomes stronger in this regime. For $B_{S}=3$ GeV2, the effect becomes even more pronounced. The peak in the intermediate region increases, while at larger momenta the ratio falls below unity. The transition point between enhancement and suppression is roughly stable across $\mu_{H}$, indicating a dynamical origin tied to the intrinsic shape of the additional function rather than to the perturbative scale choice. Overall, the results demonstrate that the inclusion of the TMDShF governed by $B_{S}$ introduces nontrivial shape modifications to the transverse-momentum distribution. Increasing $B_{S}$ amplifies these distortions, confirming its role as a strength or broadness parameter of the TMDShF. We emphasize, however, that this behavior pertains to the specific implementation described in Eq. (28), and could change under alternative NP parametrization of the CS kernel and the gluon TMDPDF.
@@
It is worth to emphasize that the present analysis has been confined to the case of electroproduction. Nonetheless, the theoretical framework developed in this work can be straightforwardly extended to provide predictions for photoproduction processes. It is, however, well established that the H1 measurements of large-$p_{T}$ $J/\psi$ photoproduction Aaron and others (2010) are accurately reproduced by the color-singlet contribution alone, computed at $\mathcal{O}(\alpha_{s}^{2})$ Flore et al. (2020), where the CO contribution at $\mathcal{O}(\alpha_{s})$ is known to significantly overshoot the experimental data. Moreover, theoretical predictions based on the conventional NRQCD factorization framework are known to fail in describing the data, particularly in the low-$p_{T}$ and high-$z$ region, as discussed in Ref. Brambilla et al. (2024a). In a preliminary analysis, as a qualitative comparison, we confront our estimate for photoproduction at $z\sim 1$ with the H1 measurement in the $0.75<z<0.9$ bin Aaron and others (2010). We find that the incorporation of the TMDShF within the TMD framework leads to theoretical predictions of the CO channel that exhibit markedly better agreement with the experimental data, yielding lower cross sections at small $p_{T}$ relative to the current collinear NRQCD results. However, we leave a detailed analysis of photoproduction in the TMD framework for future work.